OpenStudy (volleyballlover55):

Solve:

OpenStudy (volleyballlover55):

(3^8 *2^5*9^0)^-2 (2^-2/ 3^3)^4

OpenStudy (volleyballlover55):

@helpkay

OpenStudy (volleyballlover55):

show the steps it took you to solve it :)

OpenStudy (volleyballlover55):

?? @Mertsj

OpenStudy (mertsj):

\[(3^8\times2^5\times9^0)^{-2}(\frac{2^{-2}}{3^{3}})^4=\] \[\frac{1}{(3^8\times2^5)^2}\times \frac{1}{(2^2\times3^3)^4}\]

OpenStudy (volleyballlover55):

well actually the second part is divided by right?

OpenStudy (mertsj):

\[\frac{1}{3^{16}\times2^{10}}\times\frac{1}{2^8\times3^{12}}=\]

OpenStudy (mertsj):

\[\frac{1}{2^{18}\times3^{28}}\]

OpenStudy (volleyballlover55):

wait .. so whats the first step to solving this problem..? make sure to explain it in words as well please :)

OpenStudy (mertsj):

\[\frac{2^{-2}}{3^3}=\frac{1}{2^23^3}\]

OpenStudy (volleyballlover55):

first step please?? and explain it in words

OpenStudy (mertsj):

The first step was to use this property: \[a ^{-n}=\frac{1}{a^n}\]

OpenStudy (volleyballlover55):

ok and how did you use it in the problem?

OpenStudy (mertsj):

I changed 2^-2 to 1 over 2^2

OpenStudy (volleyballlover55):

look.. it would be awesome.if you can do it like this... 1. ( you explain what your doing in the step.. and then show it working it out in the actual problem.. 2.. etc..

OpenStudy (mertsj):

That would be excellent practice for you: study the steps I wrote and figure out how the laws of exponents apply and then write it down.